OpenStudy (anonymous):

2x – y = 5 3x – 2y = 12

OpenStudy (anonymous):

y=-9 x=-2

OpenStudy (anonymous):

Make the two equations equivalent to each other...in this case, instead of 2x - y = 5, you move the 5 over and get (2x - y + 5)...and instead of 3x - 2y = 12, you get (3x - 2y - 12). Now you can say (2x - y + 5) = (3x - 2y - 12). Now solve: 1. 2x - y + 5 MINUS 5 = 3x - 2y - 12 MINUS 5 2. 2x - y = 3x - 2y - 17. Now add 2y to both sides. 2a. 2x - y PLUS 2y = 3x - 2y PLUS 2y - 17 3. 2x + y = 3x - 17. Now subtract 2x from both sides. 3a. 2x MINUS 2x + y = 3x MINUS 2x - 17. 4. y = x - 17. I'm not sure if they want actual amounts for x and y, but this relationship (y is 17 greater than x) is constant. I'm assuming you're solving for the equation for a line.